OpenStudy (avanti):

lim cos 3x x aproaches pi

OpenStudy (anonymous):

you could plug it into your calcualtor

OpenStudy (anonymous):

replace \(x\) by \(\pi\)

OpenStudy (anonymous):

hopefully you do not need a calculator for \(\cos(3\pi)\)

OpenStudy (avanti):

-1?

OpenStudy (anonymous):

Correct!

OpenStudy (avanti):

ok thank you both so much!

OpenStudy (anonymous):

Your Welcome